OpenStudy (anonymous):

find dy/dx:lnsqrt(e^(3x) (x^2-1)

OpenStudy (anonymous):

(1/2(e^(3x)(x^2-1)))*(3e^(3x)*(x^2-1)+e^3x*(2x)) I think

OpenStudy (anonymous):

sorry my bad I missed something (1/2(e^(3x)(x^2-1)))*(3e^(3x)*x^2+e^(3x)*2x-3e^(3x)) = (1/(e^(3x)(2(x^2-1))))* e^(3x)(3x^2+2x-3) simplify again to get (3x^2+2x-3)/(2x^2-2) <-- final answer

OpenStudy (anonymous):

so this is the final answer?

OpenStudy (anonymous):

Yes (3x^2+2x-3)/(2x^2-2). You can check wolfram by doing d/dx lnsqrt(e^(3x) (x^2-1) = (3x^2+2x-3)/(2x^2-2), and you will see that it is true.

OpenStudy (anonymous):

cool i just confirmed thanx

OpenStudy (anonymous):

No problem do you see how I did it?

OpenStudy (anonymous):

yes i saw